OpenStudy (anonymous):

Please help Find an equation for the line: through (1, –6) and parallel to y = 2x – 2.

OpenStudy (anonymous):

parallel lines have same slope , therefore , slope of new line is 2 , the standard equation of line is y = mx + c , m is slope , c is constant plug in your values of y , x ,& m in the equation to find the constant

OpenStudy (anonymous):

get there , n I'll tell u the rest

OpenStudy (anonymous):

parallel line wud b y=2x+c passing through (1,-6) so -6=2(1)+c c=-8 so eqn is y=2x-8

OpenStudy (anonymous):

okay thank you everyone

OpenStudy (anonymous):

@RaviKiran don't just give the answer , they should understand the concept first

OpenStudy (anonymous):

check the rules and regulations

OpenStudy (anonymous):

kk i'm new here so sorry @JOELMATHEWS1234

OpenStudy (anonymous):

No probs @RaviKiran , happens to everyone